# CAJun: Continuous Adaptive Jumping using a Learned Centroidal Controller

### v1

@@ Abstract
We present CAJun, a novel hierarchical learning and control framework that enables legged robots to jump continuously with adaptive jumping distances. CAJun consists of a high-level centroidal policy and a low-level leg controller. In particular, we use reinforcement learning (RL) to train the centroidal policy, which specifies the gait timing, base velocity, and swing foot position for the leg controller. The leg controller optimizes motor commands for the swing and stance legs according to the gait timing to track the swing foot target and base velocity commands. Additionally, we reformulate the stance leg optimizer in the leg controller to speed up policy training by an order of magnitude. Our system combines the versatility of learning with the robustness of optimal control. We show that after 20 minutes of training on a single GPU, CAJun can achieve continuous, long jumps with adaptive distances on a Go1 robot with small sim-to-real gaps. Moreover, the robot can jump across gaps with a maximum width of 70cm, which is over $40\%$ wider than existing methods [^1] Videos in https://yxyang.github.io/cajun.
@@ 2 Related Works
Recently, there has been increasing interest in combining RL with optimization-based control for legged robots [22, 32, 21, 23, 33, 34]. These frameworks typically follow a hierarchical structure, where a high-level RL-trained policy outputs intermediate commands to a low-level leg controller. The RL policy can give several forms of instructions to the low-level controller, such as gait timing [22, 32], CoM trajectory [21, 34, 30, 33] and foot landing positions [23, 35, 36, 37, 38]. Our approach uses a similar hierarchical setup but adopts a general action space design where the policy specifies the gait, CoM velocity and swing foot locations simultaneously. One bottleneck of the hierarchical approaches is the slow training time because every environment rollout involves solving the optimization problem in the low-level controller. We overcome this bottleneck by relaxing the constraints in foot force optimization, so that foot force can be solved efficiently in closed form. Compared to existing frameworks which can take hours or even days to train, CAJun can be trained in 20 minutes using GPU-accelerated simulation [16].
@@ 4.2 Stance Leg Control
The stance leg controller computes the desired joint torque given the velocity command from the centroidal policy. Since jumping is mostly restricted to the sagittal plane, the policy specifies the velocity in the forward and upward axis ($v_{x},v_{z}$), as well as the rotational velocity $v_{\theta}$, and the velocity for the 3 remaining DoF is set to 00. We compute the desired torque following a 3-step procedure. First, we compute the desired CoM acceleration $\bm{\ddot{q}}^{\text{ref}}\in\mathbb{R}^{6}$ using a PD controller (Appendix. A.2). Next, we optimize for the GRF $\bm{f}=[\bm{f}_{1},\bm{f}_{2},\bm{f}_{3},\bm{f}_{4}]\in\mathbb{R}^{12}$ to track this desired acceleration, where $\bm{f}_{i}$ is the foot force vector of leg $i$. Lastly, we compute the motor torque command using $\bm{\tau}=\bm{J}^{\top}\bm{f}$, where $\bm{J}$ is the foot Jacobian. When training a hierarhical controller with a low-level optimization-based controller, a major computation bottleneck lies in the GRF optimization [21, 22, 23]. As such, we re-design this optimization procedure to significantly speed up the training process.
@@
Next, we clip the solved ground reaction forces to lie within the friction cone:
@@ 6.1 Experiment Setup
We use the Go1 quadrupedal robot from Unitree [24] as our experiment platform, and build the simulation in IsaacGym [16, 43] for fast training. To match the GPU-accelerated, massively-parallel simulation environment, we implement the entire control stack, including the centroidal policy and leg controller, in a vectorized form in PyTorch [44]. We adopt the PPO implementation from rsl_rl [16], which is specifically optimized for GPU-accelerated environments. We train CAJun on a standard desktop with an Nvidia RTX 2080Ti GPU, which takes less than 20 minutes to complete.
@@ 6.5 Comparison with End-to-End RL
To demonstrate the effectiveness of CAJun’s hierarchical setup, we compare it to an end-to-end RL baseline, where the policy directly outputs motor position commands. We use a similar MDP setup as CAJun (section. 5) for the end-to-end RL baseline. More specifically, we use the same gait generator as CAJun to generate reference foot contacts, and include stepping frequency as part of the action space so that the policy can modify the gait schedule. However, unlike CAJun, this reference gait is only used for reward computation, and does not directly affect leg controllers. We keep the same set of rewards and weights (Appendix. B.1), but add an additional alive bonus of 0.02 to ensure that it stays positive, especially during initial exploration. In both simulation and the real world, we run each policy for 6 jumps with a desired distance of 1 meter per jump, and report the total CoM displacement in Table. 2.
In the real world, CAJun outperforms the e2e baseline for both gaits (25% further in bounding, 185% in pronking). Additionally, note that CAJun faces a significantly smaller sim-to-real gap, where the total jumping distance drops by no more than 5% when transferred from sim-to-real. In comparison, the performance drop for e2e baseline can be as large as 60%. While additional sim-to-real efforts such as domain randomization [15], system identification [31] or teacher-student training [20] might reduce this gap for e2e methods, the hierarchical framework of CAJun offers a simple and efficient alternative that can be deployed zero-shot to the real world.
@@ 6.6 Ablation Study
(a) Reward Curves
(b) Training Time
Figure 7: Reward curve and training time of CAJun compared to the ablated methods.
We design a set of ablation studies to validate the design choices of CAJun. For each baseline, we report its total reward over 6 consecutive jumps with a desired distance of 1m per jump (Fig. 6(a)). We train each baseline using 5 random seeds and report the average and standard deviations. We also report the wall-clock training time in Fig. 6(b).
@@
The clipped QP in GRF optimization significantly reduced training time without noticeable performance drops. To validate this design choice, we compare the training time and policy performance of CAJun with a variant, CAJun-QP, where we solve for GRFs using the complete QP setup, where the approximated friction cone is imposed as constraints. We adopt the QP-solver from qpth [46], an efficient interior-point-method-based solver that supports GPU acceleration. For both the pronking and bounding gait, we find that CAJun achieves a similar reward compared to CAJun-QP. However, because CAJun-QP needs to iteratively optimize GRF at every control step, its training time is almost 10 times longer, which is consistent with prior observations [21].
@@ A.2 Details of the Stance Leg Controller
[Uncaptioned image] Figure 8: The centroidal dynamics model assumes that the robot is a single rigid body with massless legs. Each contact leg $i$ can generate ground reaction force $f_{i}$ at position $r_{i}$.
[Uncaptioned image] Figure 9: The reference swing trajectory is generated by interpolating between the lift-off, air and landing position. The reference trajectory is tangential to the ground, and the residuals from centroidal policy generate vertical movements.
@@
Our centroidal dynamics model is based on [8] with a few modifications (Fig. 8). We assume massless legs, and simplify the robot base to a rigid body with mass $m$ and inertia $\mathbf{I}_{\text{base}}$ (in the body frame). The rigid body dynamics in local coordinates are given by:
@@ A.3 Reference Trajectory for Swing Legs
For swing legs, we design the reference trajectory to always keep the feet tangential to the ground, and use residuals from the centroidal policy to generate vertical movements (Fig. 9). To find the reference trajectory, we interpolate between three key frames $(\bm{p}_{\text{lift-off}},\bm{p}_{\text{air}},\bm{p}_{\text{land}})$ based on the gait timing. The lift-off position $\bm{p}_{\text{lift-off}}$ is the foot location at the beginning of the swing phase. The mid-air position $\bm{p}_{\text{air}}$ is the position of the robot’s hip projected onto the ground plane. We use the Raibert Heuristic [40] to estimate the desired foot landing position:
@@
where $\bm{v}_{\text{CoM}}$ is the projected robot’s CoM velocity onto the $x-y$ plane, and $T_{\text{stance}}$ is the expected duration of the next stance phase, which is estimated using the stepping frequency from the centroidal policy. Raibert’s heuristic ensures that the stance leg will have equal forward and backward movement in the next stance phase, and is commonly used in locomotion controllers [40, 8].
@@ B.2 PPO details
As listed in Table. 3, we use the same set of hyperparameters for all PPO training, including the CAJun policies and baseline policies.

### v2

@@ Abstract
We present CAJun, a novel hierarchical learning and control framework that enables legged robots to jump continuously with adaptive jumping distances. CAJun consists of a high-level centroidal policy and a low-level leg controller. In particular, we use reinforcement learning (RL) to train the centroidal policy, which specifies the gait timing, base velocity, and swing foot position for the leg controller. The leg controller optimizes motor commands for the swing and stance legs according to the gait timing to track the swing foot target and base velocity commands. Additionally, we reformulate the stance leg optimizer in the leg controller to speed up policy training by an order of magnitude. Our system combines the versatility of learning with the robustness of optimal control. We show that after 20 minutes of training on a single GPU, CAJun can achieve continuous, long jumps with adaptive distances on a Go1 robot with small sim-to-real gaps. Moreover, the robot can jump across gaps with a maximum width of 70cm, which is over $40\%$ wider than existing methods. [^1] Video and code at this page. Author Emails: {yuxiangy,xiangyun,bboots}@cs.washington.edu, guanyas@andrew.cmu.edu, {magicmelon,tingnan,jietan}@google.com
@@ 2 Related Works
Recently, there has been increasing interest in combining RL with optimization-based control for legged robots [22, 32, 21, 23, 33, 34]. These frameworks typically follow a hierarchical structure, where a high-level RL-trained policy outputs intermediate commands to a low-level leg controller. The RL policy can give several forms of instructions to the low-level controller, such as gait timing [22, 32], CoM trajectory [21, 34, 30, 33] and foot landing positions [23, 35, 36, 37, 38]. Our approach uses a similar hierarchical setup but adopts a general action space design where the policy specifies the gait, CoM velocity and swing foot locations simultaneously. One bottleneck of the hierarchical approaches is the slow training time because every environment step involves solving the optimization problem in the low-level controller. We overcome this bottleneck by relaxing the constraints in foot force optimization [39, 40, 41], so that foot force can be solved efficiently in closed form. Compared to existing frameworks which can take hours or even days to train, CAJun can be trained in 20 minutes using GPU-accelerated simulation [16].
@@ 4.2 Stance Leg Control
The stance leg controller computes the desired joint torque given the velocity command from the centroidal policy. Since jumping is mostly restricted to the sagittal plane, the policy specifies the velocity in the forward and upward axis ($v_{x},v_{z}$), as well as the rotational velocity $v_{\theta}$, and the velocity for the 3 remaining DoF is set to 00. We compute the desired torque following a 3-step procedure. First, we compute the desired CoM acceleration $\bm{\ddot{q}}^{\text{ref}}\in\mathbb{R}^{6}$ using a PD controller (Appendix. A.2). Next, we optimize for the GRF $\bm{f}=[\bm{f}_{1},\bm{f}_{2},\bm{f}_{3},\bm{f}_{4}]\in\mathbb{R}^{12}$ to track this desired acceleration, where $\bm{f}_{i}$ is the foot force vector of leg $i$. Lastly, we compute the motor torque command using $\bm{\tau}=\bm{J}^{\top}\bm{f}$, where $\bm{J}$ is the foot Jacobian. When training a hierarchical controller with a low-level optimization-based controller, a major computation bottleneck lies in the GRF optimization [21, 22, 23]. As such, we re-design this optimization procedure to significantly speed up the training process.
@@
Next, we project the solved ground reaction forces into the friction cone, where we first clip the normal force within actuator limits, and then clip the tangential forces based on the clipped normal force:
@@
We design this projection to minimize the force disruption in the gravitational direction, so that the low-level controller can track height commands accurately.
@@ 6.1 Experiment Setup
We use the Go1 quadrupedal robot from Unitree [24], and build the simulation in IsaacGym [16, 46]. To match the GPU-accelerated simulation environment, we implement the entire control stack, including the centroidal policy and the leg controller, in a vectorized form in PyTorch [47]. We adopt the PPO implementation from rsl_rl [16]. We train CAJun on a standard desktop with an Nvidia RTX 2080Ti GPU, which takes less than 20 minutes to complete.
@@ 6.5 Comparison with End-to-End RL
To demonstrate the effectiveness of CAJun’s hierarchical setup, we compare it to an end-to-end RL baseline, where the policy directly outputs motor position commands. Please refer to Appendix. B.3 for the setup details. In both simulation and the real world, we run each policy for 6 jumps with a desired distance of 1 meter per jump, and report the total CoM displacement in Table. 2. While CAJun and end-to-end RL achieves comparable performance in simulation, CAJun faces a significantly smaller sim-to-real gap and outperforms e2e baseline for both gaits in the real world (25% further in bounding, 185% in pronking). We further conduct sim-to-sim transfer experiment and validate the robustness of CAJun under shifted dynamics (Appendix. B.3). While additional efforts such as domain randomization [15], system identification [31] or teacher-student training [20] could improve the robustness and reduce the sim-to-real gap for E2E methods, the hierarchical framework of CAJun offers a simple and efficient alternative that can be deployed zero-shot to the real world.
@@ 6.6 Ablation Study
We design a set of ablation studies to validate the design choices of CAJun. We summarize the results here. Please refer to Appendix. B.4 for details.
@@
No Swing Leg Reference
The reference swing leg trajectory improves the overall jumping performance. In NoSwingRef, we train a version of CAJun where the centroidal policy directly specify swing foot position without reference trajectory. While NoSwingRef performs similarly to CAJun for the pronking gait, it jumps significantly shorter and achieves a lower reward for the bounding gait, because the bounding gait requires more intricate coordination of swing legs.
@@
The clipped QP in GRF optimization significantly reduced training time without noticeable performance drops. To validate this design choice, we compare the training time and policy performance of CAJun with a variant, CAJun-QP, where we solve for GRFs using the complete QP setup, where the approximated friction cone is imposed as constraints. We adopt the QP-solver from qpth [49], an efficient interior-point-method-based solver that supports GPU acceleration. For both the pronking and bounding gait, we find that CAJun achieves a similar reward compared to CAJun-QP. However, because CAJun-QP needs to iteratively optimize GRF at every control step, its training time is almost 10 times longer, which is consistent with prior observations [21]. Additionally, we find that the training time speed up of CAJun can be extended to other gaits such as crawling, pacing, trotting and fly-trotting. Please refer to Appendix. B.5 for more details.
@@ 7 Limitations and Future Work
Acknowledgments
We thank He Li for helping with the motor characteristic modeling, and Philipp Wu for the design of the robot protective shell. In addition, we would like to thank Nolan Wagener, Rosario Scalise, and other friends and colleagues at the University of Washington for their support and advice through various aspects of this project.
@@ A.2 Details of the Stance Leg Controller
Our centroidal dynamics model is based on [8] with a few modifications. We assume massless legs, and simplify the robot base to a rigid body with mass $m$ and inertia $\mathbf{I}_{\text{base}}$ (in the body frame). The rigid body dynamics in local coordinates are given by:
@@ A.3 Reference Trajectory for Swing Legs
For swing legs, we design the reference trajectory to always keep the feet tangential to the ground, and use residuals from the centroidal policy to generate vertical movements. To find the reference trajectory, we interpolate between three key frames $(\bm{p}_{\text{lift-off}},\bm{p}_{\text{air}},\bm{p}_{\text{land}})$ based on the gait timing. The lift-off position $\bm{p}_{\text{lift-off}}$ is the foot location at the beginning of the swing phase. The mid-air position $\bm{p}_{\text{air}}$ is the position of the robot’s hip projected onto the ground plane. We use the Raibert Heuristic [43] to estimate the desired foot landing position:
@@
where $\bm{v}_{\text{CoM}}$ is the projected robot’s CoM velocity onto the $x-y$ plane, and $T_{\text{stance}}$ is the expected duration of the next stance phase, which is estimated using the stepping frequency from the centroidal policy. Raibert’s heuristic ensures that the stance leg will have equal forward and backward movement in the next stance phase, and is commonly used in locomotion controllers [raibert1990trotting, 8].
@@ B.1 Reward Function
B.2 PPO hyperparameters
We list the hyperparameters used in our PPO algorithm in Table. 3. We use the same set of hyperparameters for all PPO training, including the CAJun policies and baseline policies.
B.3 Comparison with End-to-End RL
E2E Setup
We use a similar MDP setup as CAJun (section. 5) for the end-to-end RL baseline. More specifically, we use the same gait generator as CAJun to generate reference foot contacts, and include stepping frequency as part of the action space so that the policy can modify the gait schedule. However, unlike CAJun, this reference gait is only used for reward computation, and does not directly affect leg controllers. For reward, we keep the same reward terms and weights (Appendix. B.1). However, since the initial exploration phase of end-to-end RL can lead to a lot of robot failures with negative rewards, we add an additional alive bonus of 0.02 to ensure that the reward stays positive.
Sim-to-Sim Transfer
To better understand the robustness of CAJun and end-to-end RL (E2E) under different dynamics, we conduct a sim-to-sim transfer experiment, where we test the performance of CAJun and E2E under increased body payloads. The result is summarized in Fig. 7. While the distance of E2E drops quickly with increased payload, CAJun maintains a near-constant distance even with a 4kg payload, thanks to the robustness of the low-level centroidal controller.
Figure 7: Comparison of total jumping distance under increased payload.
B.4 Ablation Study
Figure 8: Reward curve and jumping distance of CAJun compared to the ablated methods.
Figure 9: Training Time of CAJun compared to ablated methods.
Learning Curves
For each baseline, we report its total reward and CoM displacement over 6 consecutive jumps with a desired distance of 1m per jump (Fig. 8). We train each baseline using 5 random seeds and report the average and standard deviations. We also report the wall-clock training time in Fig. 9.
B.5 Extension to Other Gaits
While we focus on jumping gaits in this work, CAJun is a versatile locomotion framework that is capable of learning a wide range of locomotion gaits. By adopting a different contact sequence for the gait generator (Fig.2), CAJun can learn a wide variety of other locomotion gaits such as crawling, pacing, trotting and fly trotting. With GPU-parallelization, all these gaits can be trained in less than 20 minutes. Please check our website for videos.